# Towards Domain Adaptive Neural Contextual Bandits

### v1

@@ A.2 Proof for the Target Regret Bound
$\leq\epsilon_{\mathcal{T}}(h)+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
@@
$\quad+\Bigg{|}{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}-{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}\Bigg{|}$
@@
$\leq{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\epsilon_{\mathcal{T}}(h)}+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
@@
$\quad+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle+\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
@@
$\leq{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\epsilon_{\mathcal{S}}(h)+\psi+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}C_{1}}$ (13)
@@
$\quad+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\leq\epsilon_{\mathcal{S}}(h)+\psi+N\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+C_{1}+\mathit{R}_{S}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$\leq\epsilon_{\mathcal{S}}(h)+\psi+N\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+C_{1}+\mathit{R}_{S}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}+C_{2}$ (14)
$\leq\epsilon_{\mathcal{S}}(h)+\psi+N\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+\mathit{R}_{S}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}+C$ (15)
@@
$={\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}}_{Predicted~{}Reward}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}},$
where $C=C_{1}+C_{2}$. Furthermore, the inequality (13) is due to the fact that the range of reward lies in $(0,1)$, which implies $\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}})\rangle\Big{|}\bigg{]}\Bigg{)}\leq\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}1\bigg{]}\Bigg{)}:=C_{1}$, and the inequality (14) is due to $\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}})\rangle\Big{|}\bigg{]}\Bigg{)}\leq\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}})\rangle\Big{|}\bigg{]}\Bigg{)}:=C_{2}$. Furthermore, the inequality (15) is due to $\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}})\rangle\Big{|}\bigg{]}\Bigg{)}\leq\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}})\rangle\Big{|}\bigg{]}\Bigg{)}$. ∎

### v2

@@ A.2 Proof for the Target Regret Bound
$\leq$ ${\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\epsilon_{\mathcal{T}}(h)}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)}+{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\right|\right)}$
@@
$+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\right|\right)}-{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\right|\right)}$
@@
$\leq$ ${\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\epsilon_{\mathcal{T}}(h)}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)}+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\right|\right)}$
@@
$+\Bigg{|}{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\right|\right)}-{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\right|\right)}\Bigg{|}$
@@
$\leq$ ${\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\epsilon_{\mathcal{T}}(h)}+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)}$
@@
$+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle+\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)}$
[…]
@@
$+{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)}$
[…]
$+{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\mathit{R}_{S}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)}$
[…]
$+\mathit{R}_{S}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\epsilon_{\mathcal{S}}(h)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{S}}\right)}$
[…]
$+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)}_{Predicted~{}Rewards}},$
where $C=\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{T}}+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{S}}\right)$ is a small number which can be ignored as $i\to\infty$ (i.e., see Lemma A.5). ∎